Using the substitution u = 2x, solve the equation 23x = 2(2x+1) + 3(2x)


1
Expert's answer
2021-09-26T18:02:33-0400

Given equation is: 23x = 2(2x+1) + 3(2x)

Substituting 2x=u in the equation, get:

"23(\frac{u}{2})=2(u+1)+3(u)\\
\Rightarrow \frac{23u}{2}=5u+2\\
\Rightarrow \frac{23u}{2}-5u=2\\
\Rightarrow \frac{13u}{2}=2\\
\Rightarrow u=\frac{4}{13}"